OpenStudy (anonymous):

The captain of a ship at sea sights a lighthouse which is 300 feet tall. The captain measures the the angle of elevation to the top of the lighthouse to be 20 degrees. How far is the ship from the base of the lighthouse?

OpenStudy (anonymous):

First, draw the picture, set up a triangle, and find that\[\cot 20^o = \frac{x}{300} \implies x=300 \cot 20^o \]After that, you just plug it into a calculator to get the answer, or use a trig table in a book if you are a true dinosaur like me.

OpenStudy (anonymous):

What do you mean by cot?

OpenStudy (anonymous):

By the way, this solution assumes the ship captain measures his angle from the surface of the water; it's a small detail, but we have no way to really assume anything else.

OpenStudy (anonymous):

Cotangent. Are you not studying trig?

OpenStudy (anonymous):

I am but I have never heard of contangent

OpenStudy (anonymous):

It's the reciprocal of tangent.

OpenStudy (anonymous):

Tangent is opposite divided by adjacent; cotangent is adjacent divided by opposite.

OpenStudy (anonymous):

You would probably not use cotangent but let us set up the picture to see

OpenStudy (anonymous):

Could I use the inverse of tangent and plug it into my calculator

OpenStudy (anonymous):

We can tell from the picture that we have the opposite side of the angle of elevation and the adjacent side is what we are trying to find. Since we know a relationship for Opposite and Adjacent sides we can use:\[Tan \theta = Opposite/ Adjacent\]

OpenStudy (anonymous):

So since you know the relationship and have all variables except for 1 you can solve the equation for the missing variable. Do not forget to set your calculator to degrees if it is not already.

OpenStudy (anonymous):

Thank you so much I forgot to set my calculator to degrees

OpenStudy (anonymous):

My pleasure, I hope that helps!

OpenStudy (anonymous):

Not tan inverse. That gives the angle from the tan measurement. If you want to set this up using tan instead of cot, it will look like\[\tan 20^o= \frac{300}{x} \implies x=\frac{300}{\tan 20^o}\]After you figure it all out, the answers will be the same.

OpenStudy (anonymous):

Have you got this now?

OpenStudy (anonymous):

Yes thank you so much

OpenStudy (anonymous):

No sweat. Do math every day.